# LinkPrompt: Natural and Universal Adversarial Attacks on Prompt-based Language Models

### v2

@@ Abstract
Prompt-based learning is a new language model training paradigm that adapts the Pre-trained Language Models (PLMs) to downstream tasks, which revitalizes the performance benchmarks across various natural language processing (NLP) tasks. Instead of using a fixed prompt template to fine-tune the model, some research demonstrates the effectiveness of searching for the prompt via optimization. Such prompt optimization process of prompt-based learning on PLMs also gives insight into generating adversarial prompts to mislead the model, raising concerns about the adversarial vulnerability of this paradigm. Recent studies have shown that universal adversarial triggers (UATs) can be generated to alter not only the predictions of the target PLMs but also the prediction of corresponding Prompt-based Fine-tuning Models (PFMs) under the prompt-based learning paradigm. However, UATs found in previous works are often unreadable tokens or characters and can be easily distinguished from natural texts with adaptive defenses. In this work, we consider the naturalness of the UATs and develop LinkPrompt, an adversarial attack algorithm to generate UATs by a gradient-based beam search algorithm that not only effectively attacks the target PLMs and PFMs but also maintains the naturalness among the trigger tokens. Extensive results demonstrate the effectiveness of LinkPrompt, as well as the transferability of UATs generated by LinkPrompt to open-sourced Large Language Model (LLM) Llama2 and API-accessed LLM GPT-3.5-turbo.
@@ 4.6 Adaptive Defense
We compare the stealthiness of LinkPrompt with the baseline method AToP on two datasets. We select UATs generated by LinkPrompt that have comparable ASR with AToP to conduct a fair comparison. Table 6 shows the change of ASR after applying the filtering. First, we compare the drop of ASR under different trigger lengths (AToP and $\textit{LinkPrompt}_{avg}$). As shown in Table 6, the drop of ASR ($\Delta$ columns) on LinkPrompt after the filtering is overall lower than AToP on both datasets, except the result on SST-2 with trigger length 7, which indicates that LinkPrompt is more resilient to the perplexity based adaptive defense.
Second, we compare the drop of ASR under different original ASRs (indicated as $\textit{LinkPrompt}_{low}$ and $\textit{LinkPrompt}_{high}$ in Table 6), as the original low and high ASRs have a different trend. Remember we design an objective function with a weighted sum of two loss terms from the attack and the naturalness perspective respectively. We can adjust the weight $\alpha$ to control the naturalness of generated UATs. Generally, a higher $\alpha$ can result in more natural but less successful UATs, and vice versa. In Table 6, ASRs of less effective triggers ($\textit{LinkPrompt}_{low}$) even rise after the process of such a perplexity filter and the accuracy drops heavily on both tasks. This indicates the limitation of such an outlier detecting method towards LinkPrompt.
@@
$\textit{LinkPrompt}_{avg}$-3 28.72 -12.45 39.58 -19.31
$\textit{LinkPrompt}_{high}$-3 31.65 -20.73 35.15 -42.89
$\textit{LinkPrompt}_{low}$-3 25.80 -4.17 44.01 +4.27
@@
$\textit{LinkPrompt}_{avg}$-5 54.77 -20.59 48.30 -7.10
$\textit{LinkPrompt}_{high}$-5 46.07 -53.68 45.14 -27.85
$\textit{LinkPrompt}_{low}$-5 63.47 +12.51 51.46 +13.65
@@
$\textit{LinkPrompt}_{avg}$-7 63.13 +0.52 50.04 -9.71
$\textit{LinkPrompt}_{high}$-7 58.07 -15.94 50.83 -25.34
$\textit{LinkPrompt}_{low}$-7 68.18 +16.98 49.25 +5.93
@@ Acknowledgment
In this paper, we act as an attacker and present an algorithm for generating effective and natural Universal Adversarial Triggers (UATs), demonstrating strong transferability and stability as attack vectors. It is possible that the UATs or our method are being maliciously used in terms of attacking existing language models. However, researching such attacks is crucial to bolster the robustness of modern large language models. We aim to release both the algorithm and the generated triggers to aid in the development of improved defense strategies. Additionally, our experimental results contribute valuable insights into the prompt-based fine-tuning paradigm and advance our understanding of language models.
@@ B.1 Attack results of LinkPrompt on RoBERTa-large
((a)) ASR results of 3-token triggers.
((b)) ASR results of 7-token triggers.
@@ B.2 Attack results of LinkPrompt on Llama2
((a)) ASR results of 3-token triggers.
((b)) ASR results of 7-token triggers.
@@ C.1 Prompt used for naturalness measurement via ChatGPT
You are responsible for deciding which expression is more natural. Just respond with a number. Sentence1: {sen1}; Sentence2: {sen2}.
@@ C.2 Prompt used for fine-tuning Llama2
AG Predict the "[mask]" with "politics", "sports", "business" or "technology" to make the whole sentence semantically natural: Najaf battle a crucial test for Allawi Clashes between US troops and Sadr militiamen escalated Thursday, as the US surrounded Najaf for possible siege. [mask] news => politics; Galaxy, Crew Play to 0-0 Tie (AP) AP - Kevin Hartman made seven saves for Los Angeles, and Jon Busch had two saves for Columbus as the Galaxy and Crew played to a 0-0 tie Saturday night. [mask] news => sports; Wall St. Bears Claw Back Into the Black (Reuters) Reuters - Short-sellers, Wall Street’s dwindling band of ultra-cynics, are seeing green again. [mask] news => business; Oracle expands midmarket ambitions Company looks to juice its application server business with a version tuned for smaller organizations. [mask] news => technology
@@
SST2 Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: a stirring , funny and finally transporting re-imagining of beauty and the beast and 1930s horror films It was [mask]. => good; apparently reassembled from the cutting-room floor of any given daytime soap .It was [mask]. => bad
@@
IMDB Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: Not only is it a disgustingly made low-budget bad-acted movie, but the plot itself is just STUPID!!!<br /><br />A mystic man that eats women? (And by the looks, not virgin ones)<br /><br />Ridiculous!!! If you´ve got nothing better to do (like sleeping) you should watch this. Yeah right. It was [mask]. => bad; Went to see this as Me and my Lady had little else to do on a sunday afternoon I like films that deal with sleazy,loser characters and this is full of em. After a slow start we get some good turns from the cast but it is the actual ’Bellini’ that both makes and lets the film down. The ’Bellini’ is one of the funniest scenes I have seen in a film for a long while but is too short and could have made this a masterpiece overall 71/2 out of 10 It was [mask]. => good
@@
HATE Predict the "[mask]" with "harmless" or "hate" to make the whole sentence semantically natural: Happy birthday to my brother @DavonteJones10 hope you have a good day fam love you &#128170;&#127998; see you later bro [mask] speech => harmless; RT @bateson87: Send Barkley off. He’s a dirty bastard [mask] speech => hate
@@
FN Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: There was better news today for the ex-Toon hero Kevin Keegan when, after having resigned from the Newcastle United manager’s job last week, he was offered a new job: Holding Joey Barton’s Coat. Barton, still a player at United, albeit with a six-match ban, is a rabble-rouser, a trouble causer, a bit ’handy’, pushy, a ’lad’, good with his fists… temperamental, know what I mean? He regularly gets into fights, and is always in need of someone to ’ hold his coat’. The last time Barton got into a ’scuffle’, it ended in a jail sentence, and prior to that, he left Man City teammate Ousmane Dabo with his face ’caved in’. On both occasions, he was wearing a jacket, and believes he could have done so much more damage had he had a ’second’ to hold his apparel for him. Ex-boss Keegan regarded himself as something of a father figure to Barton at Newcastle, defended him in front of the Newcastle board, and stood by him when he emerged from the nick recently. Now King Kev is to support the lad permanently as he follows him around, waiting for him to explode. The job is Full Time, 24 hours a day, 7 days a week, 52 weeks a year. You get the idea. Said Keegan:" I’m lookin forward to it. He’s a nice lad, wears his heart on his sleeve, but there’s nuthin wrong with that. He’s got a fiery temperament, but he gives it his all, and you can’t ask for anythin more than that." It was [mask]. => real
@@
FR Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: This is a great product. No more fears of loosing food to bears. No assaults yet but expect it to hold up nicely. [mask] review => real; Best FPV training transition to FPV and the FPV class is a lot of fun! The other two FPV classes are a bit more complex [mask] review => fake

### v3

@@ Abstract
Prompt-based learning is a new language model training paradigm that adapts the Pre-trained Language Models (PLMs) to downstream tasks, which revitalizes the performance benchmarks across various natural language processing (NLP) tasks. Instead of using a fixed prompt template to fine-tune the model, some research demonstrates the effectiveness of searching for the prompt via optimization. Such prompt optimization process of prompt-based learning on PLMs also gives insight into generating adversarial prompts to mislead the model, raising concerns about the adversarial vulnerability of this paradigm. Recent studies have shown that universal adversarial triggers (UATs) can be generated to alter not only the predictions of the target PLMs but also the prediction of corresponding Prompt-based Fine-tuning Models (PFMs) under the prompt-based learning paradigm. However, UATs found in previous works are often unreadable tokens or characters and can be easily distinguished from natural texts with adaptive defenses. In this work, we consider the naturalness of the UATs and develop LinkPrompt, an adversarial attack algorithm to generate UATs by a gradient-based beam search algorithm that not only effectively attacks the target PLMs and PFMs but also maintains the naturalness among the trigger tokens. Extensive results demonstrate the effectiveness of LinkPrompt, as well as the transferability of UATs generated by LinkPrompt to open-sourced Large Language Model (LLM) Llama2 and API-accessed LLM GPT-3.5-turbo. The resource is available at https://github.com/SavannahXu79/LinkPrompt.
@@ 4.6 Adaptive Defense
[…]
@@
[…]
@@
[…]
@@
[…]
@@ Acknowledgement
In this paper, we introduce an algorithm for crafting powerful Universal Adversarial Triggers (UATs), showcasing their robust transferability and stability as potent attack vectors. While acknowledging potential malicious applications against language models, investigating such attacks is vital for enhancing model robustness. We intend to share both the algorithm and the generated triggers to facilitate the creation of stronger defense mechanisms. Furthermore, our experimental findings offer valuable insights into prompt-based fine-tuning and enrich our comprehension of language models.
@@ B.1 Attack results of LinkPrompt on RoBERTa-large
Refer to caption ((a)) ASR results of 3-token triggers. Refer to caption ((b)) ASR results of 7-token triggers.
@@ B.2 Attack results of LinkPrompt on Llama2
Refer to caption ((a)) ASR results of 3-token triggers. Refer to caption ((b)) ASR results of 7-token triggers.
@@ C.2 Prompt used for fine-tuning Llama2
Predict the "[mask]" with "politics", "sports", "business" or "technology" to make the whole sentence semantically natural: Najaf battle a crucial test for Allawi Clashes between US troops and Sadr militiamen escalated Thursday, as the US surrounded Najaf for possible siege. [mask] news => politics; Galaxy, Crew Play to 0-0 Tie (AP) AP - Kevin Hartman made seven saves for Los Angeles, and Jon Busch had two saves for Columbus as the Galaxy and Crew played to a 0-0 tie Saturday night. [mask] news => sports; Wall St. Bears Claw Back Into the Black (Reuters) Reuters - Short-sellers, Wall Street’s dwindling band of ultra-cynics, are seeing green again. [mask] news => business; Oracle expands midmarket ambitions Company looks to juice its application server business with a version tuned for smaller organizations. [mask] news => technology
SST2
@@
Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: a stirring , funny and finally transporting re-imagining of beauty and the beast and 1930s horror films It was [mask]. => good; apparently reassembled from the cutting-room floor of any given daytime soap .It was [mask]. => bad
IMDB
@@
Predict the "[mask]" with "bad" or "good" to make the whole sentence semantically natural: Not only is it a disgustingly made low-budget bad-acted movie, but the plot itself is just STUPID!!!<br /><br />A mystic man that eats women? (And by the looks, not virgin ones)<br /><br />Ridiculous!!! If you´ve got nothing better to do (like sleeping) you should watch this. Yeah right. It was [mask]. => bad; Went to see this as Me and my Lady had little else to do on a sunday afternoon I like films that deal with sleazy,loser characters and this is full of em. After a slow start we get some good turns from the cast but it is the actual ’Bellini’ that both makes and lets the film down. The ’Bellini’ is one of the funniest scenes I have seen in a film for a long while but is too short and could have made this a masterpiece overall 71/2 out of 10 It was [mask]. => good
HATE
@@
Predict the "[mask]" with "harmless" or "hate" to make the whole sentence semantically natural: Happy birthday to my brother @DavonteJones10 hope you have a good day fam love you &#128170;&#127998; see you later bro [mask] speech => harmless; RT @bateson87: Send Barkley off. He’s a dirty bastard [mask] speech => hate
FN
@@
Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: There was better news today for the ex-Toon hero Kevin Keegan when, after having resigned from the Newcastle United manager’s job last week, he was offered a new job: Holding Joey Barton’s Coat. Barton, still a player at United, albeit with a six-match ban, is a rabble-rouser, a trouble causer, a bit ’handy’, pushy, a ’lad’, good with his fists… temperamental, know what I mean? He regularly gets into fights, and is always in need of someone to ’ hold his coat’. The last time Barton got into a ’scuffle’, it ended in a jail sentence, and prior to that, he left Man City teammate Ousmane Dabo with his face ’caved in’. On both occasions, he was wearing a jacket, and believes he could have done so much more damage had he had a ’second’ to hold his apparel for him. Ex-boss Keegan regarded himself as something of a father figure to Barton at Newcastle, defended him in front of the Newcastle board, and stood by him when he emerged from the nick recently. Now King Kev is to support the lad permanently as he follows him around, waiting for him to explode. The job is Full Time, 24 hours a day, 7 days a week, 52 weeks a year. You get the idea. Said Keegan:" I’m lookin forward to it. He’s a nice lad, wears his heart on his sleeve, but there’s nuthin wrong with that. He’s got a fiery temperament, but he gives it his all, and you can’t ask for anythin more than that." It was [mask]. => real
FR
@@
Predict the "[mask]" with "real" or "fake" to make the whole sentence semantically natural: This is a great product. No more fears of loosing food to bears. No assaults yet but expect it to hold up nicely. [mask] review => real; Best FPV training transition to FPV and the FPV class is a lot of fun! The other two FPV classes are a bit more complex [mask] review => fake